# Beyond twisted maps: applications to motivic integration

### v2

@@ 3. Deformation theory of warps
${\mathscr{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$$\scriptstyle{\pi}$${\mathscr{T}^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\prime}}$${\operatorname{Spec}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${\operatorname{Spec}A^{\prime}}$
@@
${\mathcal{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$$\scriptstyle{\pi}$${\mathcal{Y}^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\prime}}$${\operatorname{Spec}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${\operatorname{Spec}A^{\prime}}$
@@
[…]
@@ 4.1. $\mathscr{W}\!(\mathcal{X})$ is a stack that is limit-preserving
${\mathscr{T}_{j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{j}}$$\scriptstyle{\pi_{j}}$${\mathscr{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$$\scriptstyle{\pi}$${\mathcal{X}_{A_{j}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}_{A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spec}A_{j}}$${\operatorname{Spec}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@ 4.4. Boundedness of deformations and obstructions
Propositions 4.2, 4.1, and 4.4 verify hypotheses (1)–(3), respectively, of [HR19, Main Theorem]. Proposition 4.5 shows Aff-homogeneity of $\mathscr{W}\!(\mathcal{X})$, which implies both $\textbf{Art}^{\textbf{triv}}$-homogeneity (i.e., condition (4) of (loc. cit)) and Nil-homogeneity (an implicit hypothesis in conditions (5) and (6)). Proposition 4.6 verifies conditions (5a), (5b), and (6b); conditions (5c) and (6c) are satisfied since $\operatorname{Spec}k$ is Jacobson, see condition ($\alpha$). It follows that $\mathscr{W}\!(\mathcal{X})$ is a locally finitely presented Artin stack. $\square$
@@
${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau}$${T^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathscr{W}\!(\mathcal{X})}$
@@
    ${\mathcal{X}^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{p^{\prime}}$${\mathcal{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${Y^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${Y}$
@@
${\mathscr{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{\pi}$${\mathcal{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${Y^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${Y}$
@@
${\widetilde{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{f}}$$\scriptstyle{p}$${\widetilde{Y}^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathcal{Y}^{\prime}}$
@@
which is an open subset by [GW10, Corollary 12.79]. Then $\mathcal{V}=p(\widetilde{V})$ and since $p$ is open, we see $\mathcal{V}$ is as well. ∎
@@
${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}$${\operatorname{\underline{\operatorname{Hom}}}_{T}(\mathscr{T},\mathscr{T}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$${\operatorname{\underline{\operatorname{Hom}}}_{T}(\mathscr{T},\mathcal{X}_{T})}$
@@ 7. Warped arcs with integrity
[…]
@@
to be all isomorphism classes of arcs $\gamma$ such that $\theta_{1}(\gamma)$ is $(R_{1})$-aspiring and $\theta_{0}(\gamma)$ is $(S_{1})$ and has integrity, see Definitions 9.2, 9.4, and 1.4.
@@ 9.1. Serre’s property $(S_{2})$
[…]
@@
    Let $k^{\prime\prime}/k^{\prime}$ be a finite extension of fields and let $T=\operatorname{Spec}k^{\prime}$. Then $\mathcal{D}$ is $(S_{r})$ if and only if $\mathcal{D}\times_{k^{\prime}}k^{\prime\prime}$ is.
@@
Lastly, part (3) holds since one may check $(S_{r})$tale locally and since we are working in characteristic 00, the extension $k^{\prime\prime}/k^{\prime}$ is étale. ∎
@@ 9.2. Regularity in codimension $1$
Recall from [SU21, Definition 3.1] that if $\mathcal{Y}$ is a finite type Artin stack over a field $k^{\prime}$ and $E\in D^{-}_{\operatorname{coh}}(\mathcal{Y})$, then the 00th-truncated height function $\operatorname{ht}^{(i)}_{0,E}\colon|\mathcal{Y}|\to\mathbb{Z}_{\geq 0}$ is defined by
@@
Now assume $\rho\colon V\to\mathcal{Y}$ is a finitely presented smooth cover of pure relative dimension $r$. Since $E=\Omega^{1}_{V/k^{\prime}})$ is a complex concentrated in degree 00, we see $h_{E}(v)=\dim_{K}v^{*}\Omega^{1}_{V/k^{\prime}}$ which is an upper semi-continuous function by [Sta21, tag 0BDI]. As a result, $\mathcal{Z}^{\geq d}(\Omega^{1}_{V/k^{\prime}})$ is closed. Next, from the exact triangle
@@
$0\to L^{0}y^{*}L_{\mathcal{Y}/k^{\prime}}\to v^{*}\Omega^{1}_{V/k^{\prime}}\to v^{*}\Omega^{1}_{V/\mathcal{Y}}\to L^{1}y^{*}L_{\mathcal{Y}/k^{\prime}}\to 0$
@@
Let $k^{\prime}$ be a field and $D_{1,k^{\prime}}=\operatorname{Spec}k^{\prime}[t]/t^{2}$. We say $(\mathcal{D}_{1}\to D_{1,k^{\prime}},f_{1}\colon\mathcal{D}_{1}\to\mathcal{X})\in\mathscr{L}_{1}(\mathscr{W}\!(\mathcal{X}))(k^{\prime})$ is $(R_{1})$-aspiring if $\mathcal{D}_{1}$ has $L_{\mathcal{D}_{1}/k^{\prime}}$-dimension at most $1$ in codimension 00.
@@
    If $k^{\prime\prime}/k^{\prime}$ is a finite extension of fields, then $\gamma_{1}\in\mathscr{L}_{1}(\mathscr{W}\!(\mathcal{X}))(k^{\prime})$ is $(R_{1})$-aspiring if and only if $\gamma_{1}\times_{k^{\prime}}k^{\prime\prime}$ is.
@@
For part (3), since we are working in characteristic 00, the extension $k^{\prime\prime}/k^{\prime}$ is étale. In particular, if $\gamma_{1}=(\mathcal{D}_{1,k^{\prime}}\to D_{1,k^{\prime}},\mathcal{D}_{1,k^{\prime}}\to\mathcal{X})$, then base change map $p\colon\mathcal{D}_{1,k^{\prime\prime}}=\mathcal{D}_{1,k^{\prime}}\times_{k^{\prime}}k^{\prime\prime}\to\mathcal{D}_{1,k^{\prime}}$ is codimension-preserving and $p^{*}L_{\mathcal{D}_{1,k^{\prime}}/k^{\prime}}=L_{\mathcal{D}_{1,k^{\prime\prime}}/k^{\prime\prime}}$. It follows that $\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime\prime}}/k^{\prime\prime}})=p^{-1}(\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime}}/k^{\prime}}))$, so $\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime\prime}}/k^{\prime\prime}})$ has codimension at least $1$ if and only if $(\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime}}/k^{\prime}})$ does.
@@
10. The McKay correspondence: proof of Theorems 1.4 and 1.5
@@
Using Notation 1.2, assume $Y$ is irreducible. If $f\colon\mathscr{L}(Y)\to\mathbb{Z}\cup\{\infty\}$ be a measurable function such that $\mathbb{L}^{f}$ is integrable on $\mathscr{L}(Y)$. Then
@@
Theorems 1.1, 1.2, and 1.3 tell us that the hypotheses of [SU23b, Theorem 1.2] hold. The result follows. ∎
@@
The result follows from Theorem 1.4 upon verifying the hypotheses of the theorem hold. Let $\mathcal{X}:=[V/G]$ and $Y=V/G$. We see that $V$ is of finite type and $G$ is affine, so $\mathcal{X}$ is of finite type and has affine diagonal.
Next, since $V$ is a generic $G$-representation and $\det V$ is the trivial representation (since $G\subset\operatorname{SL}(V)$), $Y$ is Gorenstein by [Kno86] (see also [SvdB17, Theorem 3.7]). As a result, [Sch05, Theorem B] implies $Y$ is log-terminal. Hence, $f:=\operatorname{ord}_{\!\mathscr{J}_{Y}}$ is measurable and $\mathbb{L}^{f}$ is integrable.
It remains to show $p\colon\mathcal{X}\to Y$ is strongly birational. By [ER21, Proposition 2.6], the stable locus $\mathcal{X}^{\operatorname{s}}\subset\mathcal{X}$ is open, its image $Y^{\operatorname{s}}\subset Y$ is open, and $\mathcal{X}^{\operatorname{s}}=\mathcal{X}\times_{Y}Y^{\operatorname{s}}$. Since there exists a point with closed orbit, $\mathcal{X}^{\operatorname{s}}$ is non-empty, hence dense. By [ER21, Proposition 2.7], $p^{\operatorname{s}}\colon\mathcal{X}^{\operatorname{s}}\to Y^{\operatorname{s}}$ factors as a gerbe over a tame stack with coarse space $Y^{\operatorname{s}}$. Specifically, the proof of [ER21, Proposition B.2] shows that the gerbe is the rigidification of $(I\mathcal{X}^{\operatorname{s}})_{0}$ in the notation of [ER21, Proposition B.6]. Since, by hypotheses, there is a stable point with trivial stabilizer, we see $(I\mathcal{X}^{\operatorname{s}})_{0}\to\mathcal{X}^{\operatorname{s}}$ is smooth with connected fibers of dimension 00, hence trivial. Thus, $\mathcal{X}^{\operatorname{s}}$ is a tame stack (which has finite inertia by definition) and $p^{\operatorname{s}}$ is a coarse space map, hence proper, quasi-finite, and a universal homeomorphism. Since $p^{\operatorname{s}}$ is proper, Theorems 2.2.5(2) and Corollary 2.2.7 of [Con07] show there is a maximal open substack $\mathcal{U}\subset\mathcal{X}^{\operatorname{s}}$ which is an algebraic space and it is characterized by the property that its geometric points have trivial automorphism groups. Since there exists a stable point with trivial stabilizer, $\mathcal{U}$ is non-empty. Since $p^{\operatorname{s}}$ is a homeomorphism, $U:=p^{\operatorname{s}}(\mathcal{U})$ is a non-empty open subset and $\mathcal{U}=U\times_{Y^{\operatorname{s}}}\mathcal{X}^{\operatorname{s}}$. It follows that $\mathcal{U}\to U$ is a coarse space map and hence an isomorphism as $\mathcal{U}$ is an algebraic space. Thus, $p$ is strongly birational. ∎

### v3

@@ 3. Deformation theory of warps
[…]
@@
[…]
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{I}\otimes(\pi_{0})_{*}\mathcal{O}_{\mathcal{Y}_{0}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi^{\prime}_{*}\mathcal{O}_{\mathcal{Y}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{*}\mathcal{O}_{\mathcal{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle$$\mathcal{I}\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\mathcal{O}_{T^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{O}_{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
@@ 4.1. $\mathscr{W}\!(\mathcal{X})$ is a stack that is limit-preserving
[…]
@@ 4.4. Boundedness of deformations and obstructions
Propositions 4.3, 4.2, and 4.5 verify hypotheses (1)–(3), respectively, of [HR19, Main Theorem]. Proposition 4.6 shows Aff-homogeneity of $\mathscr{W}\!(\mathcal{X})$, which implies both $\textbf{Art}^{\textbf{triv}}$-homogeneity (i.e., condition (4) of (loc. cit)) and Nil-homogeneity (an implicit hypothesis in conditions (5) and (6)). Proposition 4.7 verifies conditions (5a), (5b), and (6b); conditions (5c) and (6c) are satisfied since $\operatorname{Spec}k$ is Jacobson, see condition ($\alpha$). It follows that $\mathscr{W}\!(\mathcal{X})$ is a locally finitely presented Artin stack. $\square$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 7. Warped arcs with integrity
$\mathcal{Y}\xhookrightarrow{j}\mathcal{V}\xrightarrow{p}\mathcal{Y}^{\prime},$
@@
to be all isomorphism classes of arcs $\gamma$ such that $\theta_{1}(\gamma)$ is $(R_{1})$-aspiring and $\theta_{0}(\gamma)$ is $(S_{1})$ and has integrity, see Definitions 9.4, 9.7, and 1.11.
@@ 9.1. Serre’s property $(S_{2})$
    ${\mathcal{D}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{0}}$${\mathcal{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${D_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${D}$
@@
    Let $k^{\prime\prime}/k^{\prime}$ be an extension of fields and let $T=\operatorname{Spec}k^{\prime}$. Then $\mathcal{D}$ is $(S_{r})$ if and only if $\mathcal{D}\times_{k^{\prime}}k^{\prime\prime}$ is.
@@
Lastly, for part (3), we may replace $\mathcal{D}$ by a smooth cover $V\to\mathcal{D}$ to assume it is a scheme which is locally finitely presented over $k^{\prime}$. Then $V$ is $(S_{r})$ if and only if $V\times_{k^{\prime}}k^{\prime\prime}$ is. Indeed, the only if direction is given by [Sta21, Tag 0352] and the if direction follows from [Sta21, Tag 0337]. ∎
@@ 9.2. Regularity in codimension 11
Recall from [SU21, Definition 3.1] that if $\mathcal{Y}$ is a finite type Artin stack over a field $k^{\prime}$ and $E\in D^{-}_{\operatorname{coh}}(\mathcal{Y})$, then the 0th-truncated height function $\operatorname{ht}^{(i)}_{0,E}\colon|\mathcal{Y}|\to\mathbb{Z}_{\geq 0}$ is defined by
@@
Now assume $\rho\colon V\to\mathcal{Y}$ is a finitely presented smooth cover of pure relative dimension $r$. Since $E=\Omega^{1}_{V/k^{\prime}})$ is a complex concentrated in degree 0, we see $h_{E}(v)=\dim_{K}v^{*}\Omega^{1}_{V/k^{\prime}}$ which is an upper semi-continuous function by [Sta21, tag 0BDI]. As a result, $\mathcal{Z}^{\geq d}(\Omega^{1}_{V/k^{\prime}})$ is closed. Next, from the exact triangle
@@
$0\to L^{0}y^{*}L_{\mathcal{Y}/k^{\prime}}\to v^{*}\Omega^{1}_{V/k^{\prime}}\to v^{*}\Omega^{1}_{V/\mathcal{Y}}\to L^{1}y^{*}L_{\mathcal{Y}/k^{\prime}}\to 0.$
@@
Let $k^{\prime}$ be a field and $D_{1,k^{\prime}}=\operatorname{Spec}k^{\prime}[t]/t^{2}$. We say $(\mathcal{D}_{1}\to D_{1,k^{\prime}},f_{1}\colon\mathcal{D}_{1}\to\mathcal{X})\in\mathscr{L}_{1}(\mathscr{W}\!(\mathcal{X}))(k^{\prime})$ is $(R_{1})$-aspiring if $\mathcal{D}_{1}$ has $L_{\mathcal{D}_{1}/k^{\prime}}$-dimension at most $1$ in codimension 0.
@@
    If $k^{\prime\prime}/k^{\prime}$ is a field extension, then $\gamma_{1}\in\mathscr{L}_{1}(\mathscr{W}\!(\mathcal{X}))(k^{\prime})$ is $(R_{1})$-aspiring if and only if $\gamma_{1}\times_{k^{\prime}}k^{\prime\prime}$ is.
@@
For part (3), let $\gamma_{1}=(\mathcal{D}_{1,k^{\prime}}\to D_{1,k^{\prime}},\mathcal{D}_{1,k^{\prime}}\to\mathcal{X})$. Since $k^{\prime\prime}/k^{\prime}$ is flat, the base change map $p\colon\mathcal{D}_{1,k^{\prime\prime}}=\mathcal{D}_{1,k^{\prime}}\times_{k^{\prime}}k^{\prime\prime}\to\mathcal{D}_{1,k^{\prime}}$ is codimension-preserving and $p^{*}L_{\mathcal{D}_{1,k^{\prime}}/k^{\prime}}=L_{\mathcal{D}_{1,k^{\prime\prime}}/k^{\prime\prime}}$. It follows that $\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime\prime}}/k^{\prime\prime}})=p^{-1}(\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime}}/k^{\prime}}))$, so $\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime\prime}}/k^{\prime\prime}})$ has codimension at least $1$ if and only if $(\mathcal{Z}^{\geq 2}(L_{\mathcal{D}_{1,k^{\prime}}/k^{\prime}})$ does.
@@
10. The McKay correspondence: proof of Theorems 1.18 and 1.21
@@
Using Notation 1.13, assume $Y$ is irreducible. If $f\colon\mathscr{L}(Y)\to\mathbb{Z}\cup\{\infty\}$ be a measurable function such that $\mathbb{L}^{f}$ is integrable on $\mathscr{L}(Y)$. Then
@@
Theorems 1.9, 1.15, and 1.16 tell us that the hypotheses of [SU23, Theorem 1.2] hold. The result follows. ∎
The next few results aid in the proof of our McKay correspondence.
Proposition 10.2.
Let $G$ be a linearly reductive group over $k$, and let $V$ be a finite-dimensional representation of $G$.
  1. (1)
    Let $\pi:[V/G]\to[\operatorname{Spec}(k)/G]$ denote the quotient of the map $V\to\operatorname{Spec}(k)$, and let $\mathfrak{g}$ denote the adjoint representation of $G$. Then
    $\det L_{[V/G]}=\pi^{*}\left(\det V\otimes_{k}\det{\mathfrak{g}}\right).$
  2. (2)
    If there exists some $n\in\mathbb{Z}_{>0}$ such that $(\det V)^{\otimes n}$ is the trivial representation, then there exists some $m\in\mathbb{Z}_{>0}$ such that
    $(\det L_{[V/G]})^{\otimes m}\cong\mathcal{O}_{[V/G]}.$
  3. (3)
    If $\det V$ is the trivial representation and $G$ is either connected or finite, then
    $\det L_{[V/G]}\cong\mathcal{O}_{[V/G]}.$
Proof.
Let $\rho\colon V\to[V/G]=:\mathcal{X}$ be the quotient map.
From the exact triangle
$\rho^{*}L_{\mathcal{X}}\to\Omega^{1}_{V}\to\mathfrak{g}^{\vee}\otimes\mathcal{O}_{V},$
we see
$\rho^{*}(\det L_{\mathcal{X}})\simeq\det(\rho^{*}L_{\mathcal{X}})\simeq\omega_{V}\otimes\det\mathfrak{g}.$
Furthermore, as $G$-representations, $\omega_{V}$ is isomorphic to the line bundle defined by $\det V$.This therefore proves (1).
In light of (1), to prove (3) we need only show that $\det\mathfrak{g}$ is the trivial representation when $G$ is finite or connected. Since $\det\mathfrak{g}$ is the determinant of the adjoint representation, the finite case is clear and the connected case follows from [Kna96, Corollary 8.31].
Lastly, to prove (2), let $G^{0}$ be the connected component of the identity. Then the restriction of the adjoint representation of $G$ to $G^{0}$ is the adjoint representation of $G^{0}$, so it acts trivially by (3). Therefore the $G$-action on $\mathfrak{g}$ factors through $G/G^{0}$ which is finite, so we may take $m=n\cdot|G/G^{0}|$. ∎
Proposition 10.3.
Let $G$ be a linearly reductive group over $k$, and let $V$ be a generic representation of $G$. Then there exists an open subset $U\subset V/G$ such that if $\mathcal{U}$ is the preimage of $U$ along $V/G]\to V/G$, then $V/G]\smallsetminus\mathcal{U}$ has codimension at least 2 in $[V/G]$ and $\mathcal{U}\to U$ is an isomorphism.
Proof.
Let $\mathcal{X}:=[V/G]$ and $Y=V/G$. By [ER21, Proposition 2.6], the stable locus $\mathcal{X}^{\operatorname{s}}\subset\mathcal{X}$ is open, its image $Y^{\operatorname{s}}\subset Y$ is open, and $\mathcal{X}^{\operatorname{s}}=\mathcal{X}\times_{Y}Y^{\operatorname{s}}$. Since there exists a point with closed orbit, $\mathcal{X}^{\operatorname{s}}$ is non-empty, hence dense. By [ER21, Proposition 2.7], $p^{\operatorname{s}}\colon\mathcal{X}^{\operatorname{s}}\to Y^{\operatorname{s}}$ factors as a gerbe over a tame stack with coarse space $Y^{\operatorname{s}}$. Specifically, the proof of [ER21, Proposition B.2] shows that the gerbe is the rigidification of $(I\mathcal{X}^{\operatorname{s}})_{0}$ in the notation of [ER21, Proposition B.6]. Since, by hypotheses, there is a stable point with trivial stabilizer, we see $(I\mathcal{X}^{\operatorname{s}})_{0}\to\mathcal{X}^{\operatorname{s}}$ is smooth with connected fibers of dimension 0, hence trivial. Thus, $\mathcal{X}^{\operatorname{s}}$ is a tame stack (which has finite inertia by definition) and $p^{\operatorname{s}}$ is a coarse space map, hence proper, quasi-finite, and a universal homeomorphism. Since $p^{\operatorname{s}}$ is proper, Theorems 2.2.5(2) and Corollary 2.2.7 of [Con07] show there is a maximal open substack $\mathcal{U}\subset\mathcal{X}^{\operatorname{s}}$ which is an algebraic space and it is characterized by the property that its geometric points have trivial automorphism groups. Since there exists a stable point with trivial stabilizer, $\mathcal{U}$ is non-empty. Since $p^{\operatorname{s}}$ is a homeomorphism, $U:=p^{\operatorname{s}}(\mathcal{U})$ is a non-empty open subset and $\mathcal{U}=U\times_{Y^{\operatorname{s}}}\mathcal{X}^{\operatorname{s}}$. It follows that $\mathcal{U}\to U$ is a coarse space map and hence an isomorphism as $\mathcal{U}$ is an algebraic space. Finally, $V/G]\smallsetminus\mathcal{U}$ has codimension at least 2 in $[V/G]$ since $V$ is generic. ∎
Proposition 10.4.
Let $G$ be a linearly reductive group over $k$, and let $V$ be a generic representation of $G$.
  1. (1)
    If there exists some $n\in\mathbb{Z}_{>0}$ such that $(\det V)^{\otimes n}$ is the trivial representation, then $V/G$ is $\mathbb{Q}$-Gorenstein.
  2. (2)
    If $\det V$ is the trivial representation and $G$ is either connected or finite, then $V/G$ is Gorenstein.
Proof.
Set $Y=V/G$ and $\mathcal{X}=[V/G]$. We first note that $Y$ is normal by [HR74, Proposition 6.15(b)]. Now let $U\subset Y$ and $\mathcal{U}\subset\mathcal{X}$ be as in the conclusion of Proposition 10.3. We will show that $Y\smallsetminus U$ has codimension at least 2 in $Y$. Suppose to the contrary that there exists a nonempty integral subscheme $D$ of $Y\smallsetminus U$ that has codimension 1 in $Y$. Since $Y$ is normal, $D\cap Y^{\operatorname{sm}}$ is nonempty. Thus $D\cap Y^{\operatorname{sm}}$ is a nonempty Cartier divisor of $Y^{\operatorname{sm}}$, so the preimage of $D$ along $V\to V/G$ has codimension 1 in $V$, which contradicts the fact that $V/G]\smallsetminus\mathcal{U}$ has codimension at least 2 in $[V/G]$. Therefore $Y\smallsetminus U$ has codimension at least 2 in $Y$.
  1. (1)
    By Proposition 10.2(2), there exists some $m\in\mathbb{Z}_{>0}$ such that $(\det L_{\mathcal{X}})^{\otimes m}\cong\mathcal{O}_{\mathcal{X}}$. Since $U$ is isomorphic to the open substack $\mathcal{U}$ of $\mathcal{X}$, this implies that $\omega_{U}^{\otimes m}\cong\mathcal{O}_{U}$. Since $Y\smallsetminus U$ has codimension at least 2 in $Y$, this implies that $Y$ is $\mathbb{Q}$-Gorenstein.
  2. (2)
    By Proposition 10.2(3), $\det L_{\mathcal{X}}\cong\mathcal{O}_{\mathcal{X}}$. Since $U$ is isomorphic to the open substack $\mathcal{U}$ of $\mathcal{X}$, this implies that $\omega_{U}\cong\mathcal{O}_{U}$. Since $Y\smallsetminus U$ has codimension at least 2 in $Y$, this implies that $Y$ is $1$-Gorenstein. Note that $Y$ is Cohen-Macaulay by [HR74, Main Theorem], so $Y$ is Gorenstein.∎
Remark 10.5.
Proposition 10.4(2) also follows from [Kno89, Satz 2].
@@
By Proposition 10.4(1), $V/G$ is $\mathbb{Q}$-Gorenstein, so by [Sch05, Theorem B], $V/G$ is log-terminal. Therefore $\mathbb{L}^{(1/m)\operatorname{ord}_{\mathscr{J}_{V/G,m}}}$ is integrable. The theorem now follows immediately from Theorem 1.18 and Proposition 10.3. ∎